40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2194     Entry time: Fri Nov 6 16:27:15 2009
Author: Jenne 
Type: Update 
Category: PEM 
Subject: Ranger moved 

The Ranger seismometer has been moved to ~the middle of the Mode Cleaner tube, and it's orientation has been changed to horizontal (using all of the locking/mass centering procedures).  This is similar in orientation to the way things were back in the day when Rana and Matt had the OAF running nicely.

ELOG V3.1.3-